A researcher wants to estimate the average number of children with congenital heart disease who are between the ages of 1-5 years old. How many children should be enrolled in this study, if the researcher plan on using a 95% confidence level and wants a margin of error of 0.5 and standard deviation 4?


1
Expert's answer
2021-05-25T18:00:46-0400

"ME=z_{0.025}\frac{\sigma}{\sqrt{n}}."

"n=(\frac{z_{0.025}*\sigma}{ME})^2=(\frac{1.96*4}{0.5})^2=246."